# Booster: Tackling Harmful Fine-tuning for Large Language Models via Attenuating Harmful Perturbation

### v2

@@ Abstract
Harmful fine-tuning issue (Qi et al., 2023) poses serious safety concerns for Large language models’ fine-tuning-as-a-service. While existing defenses (Huang et al., 2024c; Rosati et al., 2024a) have been proposed to mitigate the issue, their performances are still far away from satisfactory, and the root cause of the problem has not been fully recovered. For the first time in the literature, we in this paper show that harmful perturbation over the model weights should be the root cause of alignment-broken of harmful fine-tuning. In order to attenuate the negative impact of harmful perturbation, we propose an alignment-stage solution, dubbed Booster. Technically, along with the original alignment loss, we append a loss regularizer in the alignment stage’s optimization. The regularizer ensures that the model’s harmful loss reduction before/after simulated harmful perturbation is attenuated, thereby mitigating the subsequent fine-tuning risk. Empirical results show that Booster can effectively reduce the harmful score of the fine-tuned models while maintaining the performance of downstream tasks. Our code is available at https://github.com/git-disl/Booster.
@@ Appendix B Baseline Descriptions
A concurrent alignment stage solution TAR, first released on August 1st (approximately one month before Booster first release) share a similar optimization viewpoint to our proposed Booster solution. We first present a simplified idea of TAR (notations is adapted for easy comparison with Booster). Specially, TAR aims to solve the following problem:
@@
where $\tilde{f}(\bm{w})$ is a representation engineering inspired retain loss for preserving performance on the capabilities proxy dataset, and $h(\bm{w})$ is the cross entropy loss over harmful data. The second term is the a tamper-resistance loss that ensures the negative cross entropy loss over the harmful data to be maximized after one step of harmful fine-tuing attack.
@@
  • $f(\bm{w})$ vs. $\tilde{f}(\bm{w})$. For optimizing the alignment loss $f(\bm{w})$, Booster need to assume the existence of alignment data. This can ensures the model learns to answer harmful prompt in a refusal way. TAR instead optimize over the retain loss, which requires the existence of a normal proxy dataset.
  • The main difference lies in the second term. TAR aims to directly maximize the harmful cross entropy loss of the model after taking the simulated harmful perturbation. In contrast, Booster aims to minimize the harmful loss reduction rate by minimizing $h(\bm{w})-h(\bm{w}-\alpha\frac{\nabla h(\bm{w})}{\|\nabla h(\bm{w})\|})$. While the difference seems to be subtle, the goal of the two designs, as we mentioned, are different. In our initial attempt, we also experiment on the format adopted by TAR, but we observe that the training would become very unstable by using TAR formulation, because the model will easily collapse ($h(\bm{w})$ will explode to a large value, and eventually leading the model to constantly repeat a single word for any prompts). Therefore, to guarantee the model’s stability, we eventually adopt the current Booster’s formulation by minimizing the harmful loss reduction rate.
@@ Appendix D Vaccine+Booster
$\min_{\bm{w}}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\max_{\|\bm{\epsilon}\|\leq\rho}\tilde{f}_{\epsilon}(\bm{w})}+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\lambda\left(h(\bm{w})-h(\bm{w}-\alpha\frac{\nabla h(\bm{w})}{\|\nabla h(\bm{w})\|})\right)}\$
where $\tilde{f}_{\epsilon}(\bm{w})$ denotes the loss over alignment dataset after applying perturbation $\epsilon$ over the embedding. The first term is the contributed from Vaccine and the second term is from Booster.
@@
$\min_{\bm{w}}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\tilde{f}_{\epsilon^{*}}(\bm{w})}+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\lambda\left(h(\bm{w})-h(\bm{w}-\alpha\frac{\nabla h(\bm{w})}{\|\nabla h(\bm{w})\|})\right)}\$

### v3

@@ Abstract
Harmful fine-tuning issue (Qi et al., 2023) poses serious safety concerns for Large language models’ fine-tuning-as-a-service. While existing defenses have been proposed to mitigate the issue, their performances are still far away from satisfactory, and the root cause of the problem has not been fully recovered. To this end, we in this paper show that harmful perturbation over the model weights could be a cause of alignment-broken of harmful fine-tuning. In order to attenuate the negative impact of harmful perturbation, we propose an alignment-stage solution, dubbed Booster. Technically, along with the original alignment loss, we append a loss regularizer in the alignment stage’s optimization. The regularizer ensures that the model’s harmful loss reduction before/after simulated harmful perturbation is attenuated, thereby mitigating the subsequent fine-tuning risk. Empirical results show that Booster can effectively reduce the harmful score of the fine-tuned models while maintaining the performance of downstream tasks. Our code is available at https://github.com/git-disl/Booster.
@@ Appendix B Baseline Descriptions
A concurrent alignment stage solution TAR (Tamirisa et al., 2024), first released on August 1st (approximately one month before Booster first release) share a similar optimization viewpoint to our proposed Booster solution. We then present a simplified idea of TAR (only one inner step in TAR is taken for sake of easy comparisons with Booster). Notations are also adapted for easy comparison with Booster. Formally, a simplified version of TAR aims to solve the following problem:
@@
where $\tilde{f}(\bm{w})$ is a representation engineering inspired retain loss for preserving performance on the capabilities proxy dataset, and $h(\bm{w})$ is the entropy loss over harmful data. The second term is a tamper-resistance loss that ensures the negative entropy loss over the harmful data to be minimized after one step of harmful fine-tuning attack. Of note, in original TAR, multiple inner steps are taken to simulate the harmful fine-tuning attack.
@@
  • $f(\bm{w})$ vs. $\tilde{f}(\bm{w})$. For optimizing the alignment loss $f(\bm{w})$, Booster need to assume the existence of alignment data. This can ensures the model learns to answer harmful prompt in a refusal way. TAR instead optimizes over the retain loss, which requires the existence of a normal proxy dataset.
  • The main difference lies in the second term. TAR aims to directly minimize the negative entropy loss over harmful data after taking the simulated harmful perturbation. In contrast, Booster aims to minimize the harmful loss reduction rate by minimizing $h(\bm{w})-h(\bm{w}-\alpha\frac{\nabla h(\bm{w})}{\|\nabla h(\bm{w})\|})$. While the difference seems to be subtle, the goal of the two designs, as we mentioned, are different. In our initial attempt, we also experiment on a term similar to tamper resistance loss used by TAR. We minimize over the negative cross entropy loss but the training would become very unstable with this loss, because the model will easily collapse ($h(\bm{w})$ will explode to a large value, and eventually leading the model to constantly repeat a single word for any prompts). Therefore, to guarantee the model’s stability, we eventually adopt the current Booster’s formulation by minimizing the harmful loss reduction rate.
@@ Appendix D Vaccine+Booster
$\min_{\bm{w}}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\max_{\|\bm{\epsilon}\|\leq\rho}\hat{f}_{\epsilon}(\bm{w})}+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\lambda\left(h(\bm{w})-h(\bm{w}-\alpha\frac{\nabla h(\bm{w})}{\|\nabla h(\bm{w})\|})\right)}\$
where $\hat{f}_{\epsilon}(\bm{w})$ denotes the loss over alignment dataset after applying perturbation $\epsilon$ over the embedding. The first term is the contributed from Vaccine and the second term is from Booster.
@@
$\min_{\bm{w}}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\hat{f}_{\epsilon^{*}}(\bm{w})}+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\lambda\left(h(\bm{w})-h(\bm{w}-\alpha\frac{\nabla h(\bm{w})}{\|\nabla h(\bm{w})\|})\right)}\$